OpenStudy (anonymous):

What are the approximate solutions of 4x^2 + 3 = −12x to the nearest hundredth?

OpenStudy (anonymous):

What are the exact solutions of x2 = 4 − 7x?

OpenStudy (anonymous):

solve \[x^2+7 x-4 = 0\] using quadratic formula get \[x=\frac{-7\pm\sqrt{65}}{21}\]

OpenStudy (anonymous):

Use complete sentences to explain how you would use graphing technology to find the solutions of 0 = −x2 + 2.5x + 1. What do the solutions of 0 = −x2 + 2.5x + 1 represent on the graph of the equation?